OpenStudy (kimjafo):

Find the roots of the quadratic equation x 2 – 8x = 9 by completing the square. Show your work.

OpenStudy (kimjafo):

I think its -1 and 9

OpenStudy (wolf1728):

x^2 -8x =9 Yes it is -1 and 9 but you need to know how to do it by completing the square.

OpenStudy (kimjafo):

I believe ive figured it out

OpenStudy (wolf1728):

x^2 -8x =9 To complete the square: Move the non-x term to the right. (It is already there). Divide the equation by the coefficient of x^2 which is one so that is already done too. Take the coefficient of 'x' , divide it by 2 then square it. -8 divided by 2 = -4 and squared = 16 Add it to both sides: x^2 -8x +16 =9 +16 Take the square root of both sides. (x-4)^2 = 5^2 x -4 = 5 Therefore x =9 and let's not forget that the square root of 25 also equals -5 so x -4 = -5 x = -1